the roots of the eqation (x-1)^3+8=0 are According to my Referance Book are -1,1-2w,1-2w^2


So please reply how is it

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear manan

earlier i have replied to this problem ,at that time u have not maintion the complex root so i was only dealing real roots.

for complex root you have to solve it

  (x-1)^3+8=0

 (x-1)3+23=0

 

you can factorise it easly with the formula

 a3 + b3 = (a+b)(a+bw)(a+bw2)    where w is the cube root of unity

 {(x-1)+2}{(x-1) +2w}{(x-1) +2w2} =0

 ( x+1)(x +2w-1)(x +2w2 -1) =0

  it gives the root  1,1-2w,1-2w^2

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin